OpenStudy (anonymous):

Can anyone tell me -15/3/5(3/5 is a fraction) Will medal and fan

OpenStudy (anonymous):

\[-15\div \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

that's the equation in case it looked confusing in the post

OpenStudy (anonymous):

When you divide by a fraction, it is the same as multiplying by the reciprocal (the fraction turned upside down). \[-15 \div\dfrac{3}{5}=-15\times\dfrac{5}{3} \]Can you tell me what the answer is now?

OpenStudy (anonymous):

okay so\[\left(\begin{matrix}-15 \\ ?\end{matrix}\right)\]

OpenStudy (anonymous):

lol thats supposed to say \[\frac{ -15 }{ 1}\]

OpenStudy (anonymous):

\[\dfrac{-15\times5}{3}=-5\times5=?\]

OpenStudy (anonymous):

then you multiply the numerators and denominators

OpenStudy (anonymous):

its supposed to be -5 or -15? :)

OpenStudy (anonymous):

-5, I divided -15 by 3 :) We could also write this as:\[\dfrac{-15\times 5}{3}=\dfrac{-75}{3} =?\]

OpenStudy (anonymous):

thats weird in my book it says keep the first fraction change division to multiplication and then replace the 2nd fraction with its reciprocal

OpenStudy (anonymous):

oh ok i see what you did you just canceled out the 1?

OpenStudy (anonymous):

oh okay so the answer is gunna be\[\frac{ -75 }{ 3}\]

OpenStudy (anonymous):

then simplify it to -25?

OpenStudy (anonymous):

Oh sorry! I see what you meant now. Do you understand this: \[-15 \div\dfrac{3}{5}=\dfrac{-15}{1}\times\dfrac{5}{3} = \dfrac{-15\times5}{1\times3}\] Yes I skipped a step as the 1 on the bottom doesn't do anything, sorry for the confusion.

OpenStudy (anonymous):

Exactly, great work!

OpenStudy (anonymous):

Thank you!